OpenStudy (anonymous):

1. Which ratios form a proportion? A. 4/9, 12/25 B. 3/7, 18/42 C. 5/11, 20/45 D. 5/8, 15/25 2. Which ratio forms a proportion with 18/30? A. 20/35 B. 50/90 C. 36/48 D. 60/100 3. Which proportion has cross products of 5× 24 and 8 ×15? A. 5/8 = 15/24 B. 5/24 = 8/15 C. 8/24 = 15/5 D. 15/8 = 5/24

OpenStudy (mathstudent55):

Start with 1. For each choice, reduce any fraction that is not reduced. If the fractions are equal, that is your answer.

OpenStudy (anonymous):

really you cant just give me the answers right away

OpenStudy (anonymous):

For example: 9:20's proportion can be 18:40.

OpenStudy (mathstudent55):

For example, choice A. Both fractions are already reduced. They are not equal, so A is not the answer.

OpenStudy (mathstudent55):

Now try choice B. 3/7, 18/42 18/42 can be reduced. Do you see a number that both 18 and 42 can be divided by?

OpenStudy (anonymous):

Now look at B..

OpenStudy (anonymous):

Exactly ^

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

So, wouldn't 6 be a factor?

OpenStudy (mathstudent55):

Yes. Divide the numerator and denominator of the right fraction of choice B by 6. \(\dfrac{18}{42} = \dfrac{18 \div 6}{42 \div 6} \) What do you get?

OpenStudy (anonymous):

Exactly, @mathstudent55

OpenStudy (anonymous):

never mind i did it myself and got a 100%

OpenStudy (anonymous):

What'd you get for 2?

OpenStudy (anonymous):

I dont understand that one that well.

OpenStudy (mathstudent55):

2. Which ratio forms a proportion with 18/30? A. 20/35 B. 50/90 C. 36/48 D. 60/100 Reduce the given fraction and reduce all the choices.

OpenStudy (mathstudent55):

2. Which ratio forms a proportion with 18/30 = 3/5? Divide by 6 A. 20/35 = 4/7 Divide by 5 B. 50/90 = 5/9 Divide by 10 C. 36/48 = 3/4 Divide by 12 D. 60/100 = 3/5 Divide by 20 Now with all fractions reduced, which fraction in the choices equals 18/30?